OpenStudy (roo777x):

what is the 40th term of the arithmetic sequence? 21 18 15 12 9 seems easy, but i forgot how to do this

OpenStudy (solomonzelman):

What pattern do you notice here?

OpenStudy (moldybubblegum11):

This is easy... in a pattern each number in the pattern is called a "term" the first number in the pattern 21 is the first term.

OpenStudy (roo777x):

all subtracted by the next variable, i get 3

OpenStudy (moldybubblegum11):

to figure this you you must figure out the pattern. to figure it out just find the difference between all the numbers.

OpenStudy (moldybubblegum11):

every other number starting with the first one gets smaller by 3. some patters will subtract and multiply, but this one just subtracts by 3.

OpenStudy (moldybubblegum11):

i mean subtract and add, sorry.

OpenStudy (moldybubblegum11):

so far you have 5 numbers in this pattern. if you want to get to the 40th term just subtract 9 by three and keep doing that 35 times.

OpenStudy (roo777x):

thanks, and i really just wanted to know how to input the formula

OpenStudy (moldybubblegum11):

or an easy way to do it is multiply 3 and 45 together because you have to just keep subtracting 3 from each term.

OpenStudy (moldybubblegum11):

3 * 45 is 105... so just subtract 105 from 9.

OpenStudy (moldybubblegum11):

the answer is -96

OpenStudy (solomonzelman):

Example: \(\color{black}{\displaystyle 3,\,7,\,11,\,15,\,17. }\) What is the \(50{\rm th}\) term of this arithmetic sequence. Solution: ((Notation: \(a_i\) is the \(i{\rm th}\) term.)) You will note that you add \(4\) to obtain the succeeding term. So, you know that \(a_1+4=a_2\) \(a_1+4+4=a_3\) \(a_1+4+4+4=a_4\) and so forth, you get, \(a_1+\underbrace{4+4+4+...+4}_{\rm (n-1)~times}=a_n\) or, alternatively, you thus may write \(a_1+\underbrace{4+4+4+...+4}_{\rm 49~times}=a_{50}\) \(a_1+4\times 49=a_{50}\) and we know that \(a_1=3,\) so, \(a_{50}=3+4\times 49=199\)

OpenStudy (moldybubblegum11):

\[\huge\color{orange}\checkmark\ \]

OpenStudy (roo777x):

by 6, i meant the next term not the 40th, whoops XD

OpenStudy (solomonzelman):

So, in general, based on the example above you will observethat for any arithmetic sequence that has a pattern of adding some number \(d\) to the succeeding term, you would have: \(\color{black}{\displaystyle a_1+(n-1)d=a_n }\).

OpenStudy (moldybubblegum11):

\[\huge\color{red}p \huge\color{orange}u \huge\color{yellow}p \huge\color{green}p \huge\color{blue}i \huge\color{magenta}e \huge\color{red}s \]

OpenStudy (solomonzelman):

Does this formula make sense to you, based on the above example?

OpenStudy (roo777x):

ok yes, thanks to both of you :{D

OpenStudy (solomonzelman):

So, if you want to find the 40th term of any sequence, in general, you then have \(\color{black}{\displaystyle a_{40}=a_1+(40-1)d }\) \(\color{black}{\displaystyle a_{40}=a_1+39d }\)

OpenStudy (solomonzelman):

So, knowing that 1. The first term is: \(\color{black}{\displaystyle a_1=21 }\) 2. The difference is: \(\color{black}{\displaystyle d=-3 }\) (u subtract 3 (or add -3) each time) you have \(\color{black}{\displaystyle a_{40}=21+39(-3) }\)

OpenStudy (roo777x):

-96

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle a_{40}=21+39(-3)=7(3)-39(3)=(7-39)(3)=(-32)(3)=-96~~~~ \checkmark }\)

OpenStudy (solomonzelman):

What I did, is just a simple demonstration of multiplying without calculator and without headaches.

OpenStudy (solomonzelman):

Anyway, the takeaway here is that: \(\color{black}{\displaystyle a_{n}=a_1+(n-1)\times d }\) where \(a_i\) denotes the \(i{\rm th}\) term, and \(d\) denotes the "pattern" that you add. This is your biggest friend for arithmetic sequence.

OpenStudy (roo777x):

thanks so much

OpenStudy (solomonzelman):

yw